OpenStudy (anonymous):

Simplify : a^4(a^-1)(a^3) a. a^7 b. a^12 c. a^6

OpenStudy (anonymous):

Use these two rules: \[\huge x^{a} \times x^{b} = x^{a + b}\]\[\huge x^{-a} = \frac{1}{x^{a}}\]

OpenStudy (anonymous):

Would it be C?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Thanks! @Calcmathlete

OpenStudy (anonymous):

np :)